RUTH GORNET

A New Construction of Isospectral Riemannian Nilmanifolds with Examples

Abstract. We present a new construction for obtaining pairs of higher-step isospectral Riemannian nilmanifolds and compare several resulting new examples. In particular, we present new examples of manifolds that are isospectral on functions, but not isospectral on one-forms.

Typeset by AMS-TEX